Consider the following reaction and answer the three questions that follow.

 2 Mg (s) + O2 (g) → 2 MgO (s)             ΔH = -1204 kJ

Calculate the amount of heat transferred when 31.67 g of Mg (s) reacts with an excess amount of oxygen gas at a constant pressure.


1
Expert's answer
2021-10-27T06:44:39-0400

MM(Mg) = 24.31 g/mol

n(Mg) "= \frac{31.67}{24.31} = 1.30 \;mol"

Energy released by 2 moles of Mg =1204 kJ

Energy released by 1.30 moles of Mg "= \frac{1204 \;kJ}{2} \times 1.3 = 782.6 \;kJ"

Answer: 782.6 kJ